How to solve this problem and please give me the answer? -8m-[2(10-5m)+15]=9m

Answer by sudhanshu_kmr(1152)


-8m-[2(10-5m)+15]=9m
=> -8m-[20-10m + 15]=9m
=> -8m-[35 - 10m}= 9m
=> -8m - 35 + 10m = 9m
=> 2 m -35 = 9m
=> 7m = -35
=> m = -35/7 = -5
so, m = -5